If a Contractor want £500 a day and you need a 17% margin, what is the charge rate to the client
If the client is paying you £610 a day, and you need to make a 18% margin, how much will you pay the contractor?

Answer by Theo(13342)

cost = 500 per day.
17% margin = 1.17 * 500 = 85 per day
charge = 585 per day.

you would have to charge at least $585 per day in order to make 17% profit.


charge = $610 per day
cost = x per day
1.18 * x = 610
solve for x to get x = $610 / 1.18 = $516.9491525 per day.

your cost would have to be less than or equal to $516.94 in order for you to make at least 18% profit.

1.18 * 516.94 = 609.982
you would make exactly 18% if you charged $609.982 per day.
since you are charging $610.00 per day, you are making a shade over 18%.